An equilateral triangle (one with all sides the same length) has an altitude of 40 inches. Find the length of the sides.
Answer by unlockmath(1599)

Hello,
One way to solve this is using the Pythagorean theorem. We set up one side as x and the other side as 1/2(x) so we have:
1/2x^2+40^2=x^2
Subtract 1/2x^2 from both sides to get:
1600=1/2x^2
Multiple by 2 to get:
3200=x^2
Sq root each side to get:
x=56.57 inches (rounded off) for each side.
Make ssense?
RJ
www.math-unlock.com